OpenStudy (winsteria):

MEDAL AND FAN Solve for x x-8 --- = -3 2 A. 2 B. –9 C. –14 D. 7

OpenStudy (anonymous):

Do *2 on both sides

OpenStudy (anonymous):

to get x-8=-6

sammixboo (sammixboo):

Do you know what to do next?

OpenStudy (winsteria):

What's next?

OpenStudy (anonymous):

Now get x by itself.

OpenStudy (anonymous):

What do you do?

sammixboo (sammixboo):

You have to add 8 on both sides of the equation to do so =) x - 8 + 8 = -6 + 8 x = -6 + 8

sammixboo (sammixboo):

So what is -6 + 8?

sammixboo (sammixboo):

So much lag :(

OpenStudy (winsteria):

2

OpenStudy (anonymous):

Yes

OpenStudy (winsteria):

So that's the answer? :)

sammixboo (sammixboo):

x = 2 =)

OpenStudy (anonymous):

A yep

OpenStudy (winsteria):

Thank youuuuuuu

OpenStudy (anonymous):

Sure :)